Previous

Calendar

May

2005

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

1

JD121
li6262 problems.

2

JD122
Site Visit (SpB): Details , Photos (29)

li6262 problems (in the morning).

from 14:00-14:45 MST, replaced two 2-micron Nupro filters for li-6262, turned off fan

east robot; done=1 (valve location=14)

3

JD123

4

JD124

5

JD125

6

JD126 (Calm/Warm)
Site Visit (SpB): Details , Photos (55)

at 9:35 MST, swiped krypton hygrometer.

All Hydra inlets flowing ok. at 14:15 MST, injected 5025ppm co2 into snowpack.

east/west robot flasks swapped (SpB)

Snow Pit (SpB)

7

JD127

8

JD128
Mothers Day

9

JD129

10

JD130

11

JD131 (Snow/Rain)

12

JD132

13

JD133
Site Visit (SpB, JH, LSD, JM, PC): Details , Photos (25)

Jeep Rental

Hydra: B3/B8 inlets have flow problems; at 14:30? MST, connected 5025ppm co2 to C4.

from 11:30-11:45 MST, Changed profiler (li-6251) dessicant

east/west robot flasks swapped (SpB/JM)

14

JD134

15

JD135

16

JD136
Hydra Inlet location soil sampled (MW)

17

JD137 (Warm/Windy)

18

JD138
Road from MRS to C1 plowed

19

JD139 (Sunny/HOT!)
Site Visit (SpB, PC): Details , Photos (41)

Compact Rental

Hydra: B3/B8 inlets slightly blocked? ; from 11-11:30 MST, adjusted span gases flow pressures .

east/west robots, done=2, reset (SpB/PC)

20

JD140

21

JD141

22

JD142

23

JD143

24

JD144
Site Visit (SpB, KR): Details , Photos (57)

from 11:00-11:30 MST, replaced domes on REBS

from 11:45-12:00 MST, turned box fans on (li-6262. li-6251, quacker)

All Hydra inlets flowing ok. at 10:15 MST, injected 5025ppm co2 into snowpack. Disconnected 5025ppm span tank.

east robot, done=0 (KR)

25

JD145
Site Visit (JH, JM): Details

"Pre-dawn" Transpiration Measurements.

26

JD146
Site Visit (SpB, KR, PC): Details , Photos (39)

Suburban Rental

from 10:00-11:30 MST, started using span gas as ref gas at Tower (instead of N2).

All Hydra inlets flowing ok. All inlets are snow-free (except for C1 and the four at profile site).

east/west robots, done=1, flasks swapped (KR/PC/SpB)

27

JD147

28

JD148
at 15:15 MST, quacker data system shuts down!! (ch200 short-haul modem dies)

29

JD149
Site Visit (SpB): Details , Photos (11)

at 10:30 MST, data system back up

Hydra data stop being saved.

30

JD150
Memorial Day

31

JD151

1

JD152
Site Visit (SpB, LSD, MW, L, KR, PC): Details , Photos (8)

Summer Jeep Rental (starts)

at 10:45 MST, chan200+hydra back on data system (start ops23)

from 11:30-11:44 MST, Changed profiler (li-6251) dessicant

Hydra data are saved again (though not switching properly?)

Soil Sampled (MW/LSD/L/PC)

2

JD153
Site Visit (JH, LSD): Details

"Pre-dawn" Transpiration Measurements.

3

JD154 (hailstorm)
Site Visit (SpB, KR): Details , Photos (32)

from 9:30-10:30 MST, using N2 as a ref gas for ec/prof systems

Hydra not working, bad relay?

east/west robots, done=1, flasks swapped (KR)

4

JD155